OpenStudy (okdutchman7):

Please help. WILL MEDAL! @jim_thompson5910 @Erak @Mehek14 @triciaal @hartnn @johnweldon1993 @satellite73

OpenStudy (triciaal):

please don't mass tag especially when the question is not posted

OpenStudy (okdutchman7):

Factor \[4x ^{2} + 8x + 28\] over the set of complex numbers.

OpenStudy (okdutchman7):

\[4(x ^{2}+2x+7)\]

OpenStudy (okdutchman7):

This is where I get stuck.

jimthompson5910 (jim_thompson5910):

4x^2+8x+28 = 4*(x^2+2x+7) now use the quadratic formula to solve x^2+2x+7 = 0

OpenStudy (okdutchman7):

I know that I have to solve it with imaginary numbers

OpenStudy (okdutchman7):

\[x = -1 \pm 6i\]

OpenStudy (okdutchman7):

\[4(x+1-6i)(x+1+6i)\]

OpenStudy (okdutchman7):

Thank you so much!

jimthompson5910 (jim_thompson5910):

incorrect. You solved x^2+2x+7 = 0 incorrectly

OpenStudy (okdutchman7):

What do you mean?

jimthompson5910 (jim_thompson5910):

you lost a root somewhere

OpenStudy (okdutchman7):

\[6i = \sqrt{-6}\]

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-2\pm \sqrt{2^2-4*1*7}}{2*1}\] \[\Large x = \frac{-2\pm \sqrt{-24}}{2}\] \[\Large x = \frac{-2\pm \sqrt{-1*4*6}}{2}\] \[\Large x = \frac{-2\pm \sqrt{-1}*\sqrt{4}*\sqrt{6}}{2}\] \[\Large x = \frac{-2\pm 2i*\sqrt{6}}{2}\] \[\Large x = -1\pm i\sqrt{6}\] `6i` is NOT the same as `i*sqrt(6)`

jimthompson5910 (jim_thompson5910):

So the full factorization would be \[\Large 4x^2+8x+28 = 4(x^2+2x+7)\] \[\Large 4x^2+8x+28 = 4(x - (-1+i\sqrt{6}))(x - (-1-i\sqrt{6}))\] \[\Large 4x^2+8x+28 = 4(x +1-i\sqrt{6})(x+1+i\sqrt{6})\]